The perimeter of a rectangular garden is 100 feet. The length of the garden is 5 feet longer than 2 times the width. What is the area of the garden ?
Answer by fractalier(6550)

Perimeter P = 2L + 2W = 100. We also have
L = 5 + 2W
Now substitute this into the above and get
2(5+2W) + 2W = 100
10 + 4W + 2W = 100
6W + 10 = 100
6W = 90
W = 15 feet
L = 5 + 2(15) = 35 feet
Area A = LW = 525 square feet